(2x + y - z = -12 3 -x + y + z = 7 1-2x + 4z = 34
Answer #1

-4-2= -7 adding 10 Qu+ y -z = -12 - Equation (il -u + y + z = 7 - Equation (ii) -2u + 4z = 34 - Equatim (iii) Now multiflyingNow : 2 + 42 -34 2 = 32-8 putting the value of u and a into Equatim (ii) - ty tz = 7 - (4) +y + 8 = 7 on y + 9 = 7 y=-2 V=-1,